# Quantum Mechanical Softening of the Hypertriton Transverse Momentum Spectrum in Heavy-Ion Collisions

### v2

@@ Abstract
Understanding the properties of hypernuclei helps to constrain the interaction between hyperon and nucleon, which is known to play an essential role in determining the properties of neutron stars. Experimental measurements have suggested that the hypertriton (${}^{3}_{\Lambda}\text{H}$), the lightest hypernucleus, exhibits a halo structure with a deuteron core encircled by a $\Lambda$ hyperon at a distance of about 10 fm. This large $\Lambda-d$ distance in ${}^{3}_{\Lambda}\text{H}$ wave function is found to cause a suppressed ${}^{3}_{\Lambda}\text{H}$ yield and a softening of its transverse momentum ($p_{T}$) spectrum in relativistic heavy-ion collisions. Within the coalescence model based on nucleons and $\Lambda$ hyperons from a microscopic hybrid hydro model with a hadronic afterburner for nuclear cluster production in Pb-Pb collisions at $\sqrt{s_{NN}}$= 5.02 TeV, we show how this softening of the hypertriton $p_{T}$ spectrum appears and leads to a significantly smaller mean $p_{T}$ for ${}^{3}_{\Lambda}\text{H}$ than for helium-3 (3He). The latter is opposite to the predictions from the blast-wave model which assumes that ${}^{3}_{\Lambda}\text{H}$ and 3He are thermally produced at the kinetic freeze-out of heavy ion collisions. The discovered quantum mechanical softening of the (anti-)hypertriton spectrum can be experimentally tested in relativistic heavy-ion collisions at different collision energies and centralities and used to obtain valuable insights to the mechanisms for light (hyper-)nuclei production in these collisions.
@@
Introduction. Observations of anti-nuclei such as anti-deuteron ($\bar{d}$), anti-helium (${}^{3}\overline{\text{He}}$ and ${}^{4}\overline{\text{He}}$), and anti-hypertriton (${}^{3}\overline{{}_{\bar{\Lambda}}\text{H}}$) have been reported in many heavy-ion collision experiments Abelev et al. (2010); Agakishiev et al. (2011); Adam et al. (2015a, 2020); Chen et al. (2018); Braun-Munzinger and Dönigus (2019); Acharya et al. (2023a). Recent experiments at the Relativistic Heavy Ion Collider (RHIC) Aboona et al. (2023a) and the Large Hadron Collider (LHC) ALI (2023) have even suggested possible detection of heavier anti-hypernuclei, such as $\overline{{}^{4}_{\bar{\Lambda}}\text{H}}$ and $\overline{{}^{4}_{\bar{\Lambda}}\text{He}}$. The detection of these bound nuclei and the study of their properties Gal et al. (2016); Adamczyk et al. (2015); Acharya et al. (2023b, 2019); Aboona et al. (2023b) are important in the search for the signals of phase transitions in the strongly interacting matter created in relativistic heavy-ion collisions Sun et al. (2017, 2018, 2021); Shuryak and Torres-Rincon (2019, 2020); Abdulhamid et al. (2023); Ko (2023) and in the dark matter detection in space Blum et al. (2017); von Doetinchem et al. (2020); Acharya et al. (2023c) as well as for understanding the fundamental CPT theorem in quantum field theory Adam et al. (2015a, 2020). Also, understanding the properties of hypernuclei helps to constrain the hyperon-nucleon ($Y\text{-}N$) interaction Chen et al. (2023); Ma (2023); Zhang et al. (2022), which is essential for determining the structure of compact astrophysical objects like the neutron stars Gal et al. (2016).
@@
In the coalescence model Scheibl and Heinz (1999); Sun et al. (2019) for light cluster production, the formation probability of a hypertriton from a proton, a neutron, and a $\Lambda$ hyperon is given by the Wigner function of the hypertriton internal wave function, which we take as
$W_{\text{ht}}$ $=$ $8^{2}g_{\text{ht}}e^{-\frac{\rho^{2}}{\sigma_{1}^{2}}-\sigma_{1}^{2}k_{\rho}^{2}}e^{-\frac{\lambda^{2}}{\sigma_{2}^{2}}-\sigma_{2}^{2}k_{\lambda}^{2}},$ (1)
where $g_{ht}=1/4$ is the statistical factor for spin 1/2 proton, neutron, and $\Lambda$ hyperon to form a spin 1/2 hypertriton. The relative coordinates and momenta in the above equation are defined by
@@
$k_{\lambda}$ $=$ $\sqrt{\frac{3}{2}}\frac{m_{3}({\bf r}_{1}+{\bf r}_{2})-(m_{1}+m_{2}){\bf r}_{3}}{m_{1}+m_{2}+m_{3}}.$ (4)
Here, ${\bf r}_{1}$, ${\bf r}_{2}$, and ${\bf r}_{3}$ are the spatial coordinates of neutron, proton, and $\Lambda$ hyperon, respectively at an equal time in their rest frame, and they are determined by propagating the constituent particles at earlier freeze-out time to the time of the last freeze-out one. The size parameter $\sigma_{1}$ in the Wigner function is related to the deuteron root-mean-squared radius $r_{{d}}$ by $\sigma_{1}=\sqrt{4/3}~{}r_{d}\approx 2.26$ fm Sun and Chen (2017); Ropke (2009). The size parameter $\sigma_{2}$ is related to the $\Lambda-d$ distance by $\sigma_{2}=\frac{2}{3}\sqrt{\langle r_{\Lambda d}^{2}\rangle}$, and it can be parameterized as $\sigma_{2}\approx(2.15(B_{\Lambda}/\text{MeV})^{-1/2}+1.23)~{}$fm, with details provided in the AppendixA. For the binding energies measured by the STAR and ALICE collaboration, we obtain $\sigma_{2}=5.45$ fm and $\sigma_{2}=7.96$ fm, respectively, while using the world averaged $B_{\Lambda}$ value gives $\sigma_{2}=6.52$ fm. For helium-3 production in the coalescence model, we use $\sigma_{1}=\sigma_{2}\approx 1.76$ fm Ropke (2009) as it has a spherical shape.
@@
Figure 1 displays the centrality dependence of the yield ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$  for different values of $B_{\Lambda}$. The yield ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$  is seen to be suppressed in more peripheral collisions (30-50%) as compared to central collisions (0-10%). This suppression is due to the decrease of the emission source volume with increasing collision centrality. Also, the yield ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$  is suppressed for smaller $\Lambda$-separation energies or larger $\Lambda-d$ distance. The suppression of hypertriton production in non-central collisions is in accordance with earlier studies on light cluster production in collisions of small systems like $pp$ or $p$+Pb collisions at the LHC energies Sun et al. (2019).
@@
Softening of hypertriton $p_{T}$ spectrum Figure 2 displays by shaded bands the ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ as a function of transverse momentum ($p_{T}$) from the coalescence model. It is observed that for all three different values of $\Lambda$ separation energies and all centralities, there is a pronounced decreasing trend in this ratio with an increase in $p_{T}$. This trend is in sharp contrast with the behavior predicted by the blast-wave (BW) model, shown by solid lines, which shows an increase of ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ with increasing $p_{T}$. In the BW model, both hypertriton and helium-3 are assumed in local thermal equilibrium at their kinetic freeze-out with an invariant momentum spectrum $f=E\text{d}^{3}N/\text{d}p^{3}$ given by Cooper and Frye (1974b); Schnedermann et al. (1993):
@@
In the above, $I_{0}$ and $K_{1}$ are the modified Bessel functions with their arguments depending on the transverse mass $m_{T}=\sqrt{m^{2}+p_{T}^{2}}$, the kinetic freeze-out temperature $T_{\rm kin}$, and the flow rapidity profile $\rho=\tanh^{-1}[(r/R_{0})^{n}\beta_{s}]$ with $r$ and $R_{0}$, $\beta_{s}$, and $n$ being the radial distance in the transverse plane, the transverse radius, the transverse expansion velocity at the surface, and the exponent of the velocity profile, respectively. The blast-wave model can be viewed as a simple approximation to the hydrodynamic model used in describing heavy ion collisions. The results shown in Fig. 2 for the blast-wave model are obtained by using the fitted kinetic freeze-out temperature ($T_{\text{kin}}$), flow velocity ($\beta_{s}$) and exponent ($n$) in Ref. Acharya et al. (2020c), i.e., 0.090 GeV, 0.907, and 0.735 for 0-10% centrality, 0.094 GeV, 0.90, and 0.739 for 10-30% centrality, and 0.105 GeV, 0.88, and 0.828 for 30-50% centrality. For large $m_{T}$, $f_{\text{BL}}$ takes the form $f_{\text{BL}}\propto\text{exp}(-m_{T}/T_{\text{eff}})$ with $T_{\text{eff}}=T_{\rm kin}\sqrt{\frac{1+\langle\beta_{T}\rangle}{1-\langle\beta_{T}\rangle}}$ being the effective (blue-shifted) temperature. The ratio of the hypertriton to the helium-3 spectrum at large $m_{T}$ is then given by
@@
which increases as a function of $p_{T}$ because the transverse flow drives the heavier hypertriton to higher momentum than the lighter helium-3, and this explains the increasing trend of the blast-wave model results (solid lines) shown in Fig. 2.
Figure 3: Ratio $\langle p_{T}\rangle_{{}^{3}_{\Lambda}\text{H}}/\langle p_{T}\rangle_{{}^{3}\text{He}}$ of hypertriton to helium-3 as a function of $\Lambda-d$ distance in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV for 0-10% (a), 10-30% (b), and 30-50% (c) centralities. The solid lines denote the prediction of the blast-wave model, whereas the shaded bands denote the prediction from the coalescence model. The solid symbols correspond to results obtained for various values of $\Lambda$ separation $B_{\Lambda}$.
Suppression of hypertriton mean transverse momentum Figure 3 further shows the ratio ($\langle p_{T}\rangle_{{}^{3}_{\Lambda}\text{H}}/\langle p_{T}\rangle_{{}^{3}\text{He}}$) of hypertriton to helium-3 mean-$p_{T}$ as a function of $\Lambda-d$ distance. The solid lines denote the prediction of the blast-wave model, whereas the shaded bands denote the prediction from the coalescence model. The solid symbols correspond to results obtained with various values of $B_{\Lambda}$. According to the blast-wave model prediction, the value of the hypertriton to helium-3 mean-$p_{T}$ ratio has a slightly larger than unity value of about 1.03-1.04. The coalescence model predicts, on the other hand, a value for this ratio that is substantially smaller than one for all three values of $B_{\Lambda}$, and it decreases with increasing $\Lambda-d$ distance or decreasing $B_{\Lambda}$. This trend of a smaller mean $p_{T}$ for hypertriton than that for helium-3 is seen for all collision centralities. The centrality dependence of the softening effect shown in Fig. 3 is, however, rather weak, and this may be attributed to the reason discussed qualitatively below.
@@
Within the coalescence model, the ratio of hypertriton spectrum to helium-3 spectrum is approximately given by Scheibl and Heinz (1999); Bellini and Kalweit (2019); Blum and Takimoto (2019)
$\frac{f_{\text{COAL}}^{\text{ht}}}{f_{\text{COAL}}^{\text{he3}}}\approx\frac{f_{\text{BL}}^{\text{ht}}}{f_{\text{BL}}^{\text{he3}}}\times\frac{[1+\frac{\sigma_{\text{he3}}^{2}}{2R^{2}(p_{T})}]^{3}}{[1+\frac{\sigma_{1}^{2}}{2R^{2}(p_{T})}]^{3/2}[1+\frac{\sigma_{2}^{2}}{2R^{2}(p_{T})}]^{3/2}},$ (7)
@@
where the last term (denoted by $\mathcal{C}$ hereafter) in the above equation is due to the quantum mechanical corrections Scheibl and Heinz (1999) from the hypertriton wave function. Because the inhomogeneity size $R(p_{T})$ decreases with increasing $p_{T}$ Adam et al. (2015b) and $\sigma_{2}\gg\sigma_{\text{he3}}\sim\sigma_{1}$, the quantum correction factor $\mathcal{C}$ leads to a suppression the ${}^{3}_{\Lambda}\text{H}$ yield at larger $p_{T}$. For a large emission source with $R\geq\sigma_{2}$, the correction factor can be expressed as
$\mathcal{C}\approx 1-\frac{3(\sigma_{2}^{2}-\sigma_{1}^{2}-\sigma_{\text{he3}}^{2})}{2R(p_{T})^{2}},$ (8)
@@
which becomes less than one when $\sigma_{2}\gtrsim 4$ fm, resulting in a large quantum correction and thus a decreasing ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ ratio with increasing $p_{T}$ shown in Fig. 2 at all collision centralities. The $p_{T}$ dependence of the quantum correction factor $\mathcal{C}$ has a different behavior for a small emission source with $R\leq\sigma_{\text{he3}}$. In this case, it can be simplified as
$\mathcal{C}\approx\frac{\sigma_{\text{he3}}^{6}}{\sigma_{1}^{6}\sigma_{2}^{6}}\times\left[1-6R^{2}(p_{T})\left(\frac{1}{\sigma_{\text{he3}}^{2}}-\frac{1}{\sigma_{1}^{2}}-\frac{1}{\sigma_{2}^{2}}\right)\right],$ (9)
which shows a weaker $p_{T}$ dependence than in the case of a large emission source because $\sigma_{\text{he3}}\sim\sigma_{1}$, making the last term in the above equation close to zero. This effect due to the relative hypertrition and source sizes indicates that the softening of hypertriton transverse momentum is weaker in collisions of small systems like $pp$ and $pA$ collisions.
@@
Conclusion and outlook. In the present study, we have investigated hypertriton production in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV using the coalescence model with kinetic freeze-out nucleons and $\Lambda$ hyperons from a microscopic hybrid approach based on the MUSIC hydrodynamic model and the UrQMD hadronic transport model. We have found that the halo structure of hypertriton with a large $\Lambda-d$ distance of approximately $10$ fm leads to not only a suppression of ${}^{3}_{\Lambda}\text{H}$ yield but also a softening of its transverse momentum ($p_{T}$) spectrum with a weak centrality dependence. In particular, the mean $p_{T}$ of ${}^{3}_{\Lambda}\text{H}$ is found to be significantly smaller than that of helium-3 even in the most central collisions, which is in sharp contrast with the predictions of the blast-wave model. Such a quantum mechanical softening on (anti-)hypertriton spectrum is a general feature and a natural outcome of the hypertriton wave function used in the final-state coalescence model, which can be readily tested in high-energy experiments with different beam energies and collision systems, providing thus a good opportunity to unravel the production mechanism of (anti-)hypernuclei in high-energy nuclear collisions and also to obtain constraints on the $\Lambda$-nucleon interaction.

### v4

# Softening of the Hypertriton Transverse Momentum Spectrum in Heavy-Ion Collisions

@@ Abstract
Understanding the properties of hypernuclei helps to constrain the interaction between hyperon and nucleon, which is known to play an essential role in determining the properties of neutron stars. Experimental measurements have suggested that the hypertriton (${}^{3}_{\Lambda}\text{H}$), the lightest hypernucleus, exhibits a halo structure with a deuteron core encircled by a $\Lambda$ hyperon at a distance of about 10 fm. This large $\Lambda-d$ distance in ${}^{3}_{\Lambda}\text{H}$ wave function is found to cause a suppressed ${}^{3}_{\Lambda}\text{H}$ yield and a softening of its transverse momentum ($p_{T}$) spectrum in relativistic heavy-ion collisions. Within the coalescence model based on nucleons and $\Lambda$ hyperons from a microscopic hybrid hydro model with a hadronic afterburner for nuclear cluster production in Pb-Pb collisions at $\sqrt{s_{NN}}$= 5.02 TeV, we show how this softening of the hypertriton $p_{T}$ spectrum appears and leads to a smaller mean $p_{T}$ for ${}^{3}_{\Lambda}\text{H}$ than for helium-3 (3He). The latter is opposite to the predictions from the blast-wave model which assumes that ${}^{3}_{\Lambda}\text{H}$ and 3He are thermally produced at the kinetic freeze-out of heavy-ion collisions. The discovered quantum mechanical softening of the (anti-)hypertriton spectrum can be experimentally tested in relativistic heavy-ion collisions at different collision energies and centralities and used to obtain valuable insights to the mechanisms for light (hyper-)nuclei production in these collisions.
@@
Introduction. Observations of anti-nuclei such as anti-deuteron ($\bar{d}$), anti-helium (${}^{3}\overline{\text{He}}$ and ${}^{4}\overline{\text{He}}$), and anti-hypertriton (${}^{3}_{\bar{\Lambda}}\overline{\text{H}}$) have been reported in many heavy-ion collision experiments Abelev et al. (2010); Agakishiev et al. (2011); Adam et al. (2015a, 2020); Chen et al. (2018); Braun-Munzinger and Dönigus (2019); Acharya et al. (2023a). Recent experiments at the Relativistic Heavy Ion Collider (RHIC) Aboona et al. (2023a) and the Large Hadron Collider (LHC) ALI (2023) have even suggested possible detection of heavier anti-hypernuclei, such as ${}^{4}_{\bar{\Lambda}}\overline{\text{H}}$ and ${}^{4}_{\bar{\Lambda}}\overline{\text{He}}$. The detection of these bound nuclei and the study of their properties Gal et al. (2016); Adamczyk et al. (2015); Acharya et al. (2023b, 2019); Aboona et al. (2023b) are important in the search for the signals of phase transitions in the strongly interacting matter created in relativistic heavy-ion collisions Sun et al. (2017, 2018, 2021); Shuryak and Torres-Rincon (2019, 2020); Abdulhamid et al. (2023); Ko (2023) and in the dark matter detection in space Blum et al. (2017); von Doetinchem et al. (2020); Acharya et al. (2023c) as well as for understanding the fundamental CPT theorem in quantum field theory Adam et al. (2015a, 2020). Also, understanding the properties of hypernuclei helps to constrain the hyperon-nucleon ($Y\text{-}N$) interaction Chen et al. (2023); Ma (2023); Zhang et al. (2022), which is essential for determining the structure of compact astrophysical objects like the neutron stars Gal et al. (2016).
@@
In the coalescence model Scheibl and Heinz (1999); Sun et al. (2019) for light cluster production, the formation probability of a hypertriton from a proton, a neutron, and a $\Lambda$ hyperon is given by the product of the statistical factor $g_{ht}=1/4$ for spin 1/2 proton, neutron, and $\Lambda$ hyperon to form a spin 1/2 hypertriton and the Wigner function of the hypertriton internal wave function, which we take as
$W_{\text{ht}}$ $=$ $8^{2}e^{-\frac{\rho^{2}}{\sigma_{1}^{2}}-\sigma_{1}^{2}k_{\rho}^{2}}e^{-\frac{\lambda^{2}}{\sigma_{2}^{2}}-\sigma_{2}^{2}k_{\lambda}^{2}}.$ (1)
The relative coordinates and momenta in the above equation are defined by
@@
$k_{\lambda}$ $=$ $\sqrt{\frac{3}{2}}\frac{m_{3}({\bf k}_{1}+{\bf k}_{2})-(m_{1}+m_{2}){\bf k}_{3}}{m_{1}+m_{2}+m_{3}}.$ (4)
Here, ${\bf r}_{1}$, ${\bf r}_{2}$, and ${\bf r}_{3}$ are the spatial coordinates of neutron, proton, and $\Lambda$ hyperon, respectively, at an equal time in their rest frame, and they are determined by propagating the constituent particles of momenta ${\bf k}_{1}$, ${\bf k}_{2}$, and ${\bf k}_{3}$ at earlier freeze-out times to the time of the last freeze-out one. The size parameter $\sigma_{1}$ in the Wigner function is related to the deuteron root-mean-squared radius $r_{{d}}$ by $\sigma_{1}=\sqrt{4/3}~{}r_{d}\approx 2.26$ fm Sun and Chen (2017); Ropke (2009). The size parameter $\sigma_{2}$ is related to the $\Lambda-d$ distance by $\sigma_{2}=\frac{2}{3}\sqrt{\langle r_{\Lambda d}^{2}\rangle}$, and it can be parameterized as $\sigma_{2}\approx(2.15(B_{\Lambda}/\text{MeV})^{-1/2}+1.23)~{}$fm, with details provided in the AppendixA. For the binding energies measured by the STAR and ALICE collaboration, we obtain $\sigma_{2}=5.45$ fm and $\sigma_{2}=7.96$ fm, respectively, while using the world averaged $B_{\Lambda}$ value gives $\sigma_{2}=6.52$ fm. For helium-3 production in the coalescence model, we use $\sigma_{1}=\sigma_{2}\approx 1.76$ fm Ropke (2009) as it has a spherical shape.
@@
We note that the statistical factor $g_{\rm ht}=1/4$ used in the coalescence model for proton, neutron, and $\Lambda$ hyperon to form a hypertriton does not include that due to their isospins. As shown in Ref. Sun et al. (2018), the light nuclei numbers from such a coalescence model based on nucleons from a thermal source reduce to those from the thermal model when their binding energies are much smaller than the temperature and their radii are much larger than the size of the particle emission source. The exclusion of the isospins in the coalescence model based on kinetically freeze-out nucleons in relativistic heavy-ion collisions is also supported by the results in Ref. Sun et al. (2024) from the kinetic approach, which treats light nuclei as dynamic degrees of freedom by including their production and annihilation from pion catalyzed multi-nucleon reactions during the expansion stage of heavy-ion collisions, because similar light nuclei numbers are obtained from these two approaches.
Figure 1 displays the collision centrality dependence of the yield ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$  for different values of $B_{\Lambda}$. The three collision centralities of 0-10%, 10-30%, and 30-50% in this figure correspond to the charged-particle multiplicity $dN_{ch}/d\eta$ of about 1700, 1040, and 440, respectively, which are consistent with those measured in ALICE experiments Acharya et al. (2020c). The yield ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$  is seen to be suppressed in more peripheral collisions (30-50%) as compared to central collisions because of increasing emission source volume with decreasing collision centrality. Also, the yield ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$  is more suppressed for a smaller $\Lambda$-separation energy or a larger $\Lambda-d$ distance with a variation by a factor of 2-3 for the different $\Lambda$-separation energies considered in the present study. The suppression of hypertriton production in non-central collisions is in accordance with earlier studies on light cluster production in collisions of small systems like $pp$ or $p$+Pb collisions at the LHC energies Sun et al. (2019).
Figure 3: Ratio $\langle p_{T}\rangle_{{}^{3}_{\Lambda}\text{H}}/\langle p_{T}\rangle_{{}^{3}\text{He}}$ of hypertriton to helium-3 as a function of $\Lambda-d$ distance in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV for 0-10% (a), 10-30% (b), and 30-50% (c) centralities. The solid lines denote the prediction of the blast-wave model, whereas the shaded bands denote the prediction from the coalescence model. The solid symbols correspond to results obtained for various values of $\Lambda$ separation energy $B_{\Lambda}$.
Softening of hypertriton $p_{T}$ spectrum Figure 2 displays by shaded bands the ratio ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ as a function of transverse momentum ($p_{T}$) from the coalescence model. It is observed that this spectrum ratio in all centralities also varies by a factor of 2-3 among the three different values of the $\Lambda$ separation energy as for the yield ratio shown in Fig. 1. Also, there is a pronounced decreasing trend in this ratio with an increase in $p_{T}$. This trend is in sharp contrast with the behavior predicted by the blast-wave (BW) model, shown by solid lines, which shows an increase of ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ with increasing $p_{T}$. In the BW model, both hypertriton and helium-3 are assumed in local thermal equilibrium at their kinetic freeze-out with an invariant momentum spectrum $f=E\text{d}^{3}N/\text{d}p^{3}$ given by Cooper and Frye (1974b); Schnedermann et al. (1993):
@@
In the above, $I_{0}$ and $K_{1}$ are the modified Bessel functions with their arguments depending on the transverse mass $m_{T}=\sqrt{m^{2}+p_{T}^{2}}$, the kinetic freeze-out temperature $T_{\rm kin}$, and the flow rapidity profile $\rho=\tanh^{-1}[(r/R_{0})^{n}\beta_{s}]$ with $r$ and $R_{0}$, $\beta_{s}$, and $n$ being the radial distance in the transverse plane, the transverse radius, the transverse expansion velocity at the surface, and the exponent of the velocity profile, respectively. The blast-wave model can be viewed as a simple approximation to the hydrodynamic model used in describing heavy-ion collisions. The results shown in Fig. 2 for the blast-wave model are obtained by using the fitted kinetic freeze-out temperature ($T_{\text{kin}}$), flow velocity ($\beta_{s}$) and exponent ($n$) in Ref. Acharya et al. (2020c), i.e., 0.090 GeV, 0.907, and 0.735 for 0-10% centrality, 0.094 GeV, 0.90, and 0.739 for 10-30% centrality, and 0.105 GeV, 0.88, and 0.828 for 30-50% centrality. For large $m_{T}$, $f_{\text{BL}}$ takes the form $f_{\text{BL}}\propto\text{exp}(-m_{T}/T_{\text{eff}})$ with $T_{\text{eff}}=T_{\rm kin}\sqrt{\frac{1+\langle\beta_{T}\rangle}{1-\langle\beta_{T}\rangle}}$ being the effective (blue-shifted) temperature. The ratio of the hypertriton to the helium-3 spectrum at large $m_{T}$ is then given by
@@
which increases with increasing $p_{T}$ because the transverse flow drives the heavier hypertriton to higher momentum than the lighter helium-3, and this explains the increasing trend of the blast-wave model results (solid lines) shown in Fig. 2.
@@
In the coalescence model, the ratio of hypertriton to helium-3 $p_{T}$ spectra is, on the other hand, approximately given by Scheibl and Heinz (1999); Bellini and Kalweit (2019); Blum and Takimoto (2019)
$\frac{f_{\text{COAL}}^{\text{ht}}}{f_{\text{COAL}}^{\text{he3}}}\approx\frac{f_{\text{BL}}^{\text{ht}}}{f_{\text{BL}}^{\text{he3}}}\frac{[1+\frac{\sigma_{\text{he3}}^{2}}{2R^{2}(p_{T})}]^{3}}{[1+\frac{\sigma_{1}^{2}}{2R^{2}(p_{T})}]^{3/2}[1+\frac{\sigma_{2}^{2}}{2R^{2}(p_{T})}]^{3/2}},$ (7)
@@
where $R(p_{T})$ is the inhomogeneity Gaussian size of the nucleon and Lambda emission source for produced hypertrion and helium-3 of transverse momentum $p_{T}$, and the last factor (denoted by $\mathcal{C}$ hereafter) is from the quantum mechanical corrections Scheibl and Heinz (1999); Sun et al. (2019) due to the hypertriton wave function. Because of the decreasing $R(p_{T})$ with increasing $p_{T}$ Adam et al. (2015b) and the relation $\sigma_{2}\gg\sigma_{\text{he3}}\sim\sigma_{1}$, the quantum correction factor $\mathcal{C}$ leads to a suppression of the ${}^{3}_{\Lambda}\text{H}$ yield at larger $p_{T}$ compared to that of helium-3. This can be explicitly seen as follows. Since $R(p_{T})$ in the collisions considered in our study is larger than the sizes of hypertriton and helium-3, i.e., $R(p_{T})\geq\sigma_{2}\gg\sigma_{1},\sigma_{\rm he3}$, the correction factor can be approximately expressed as
$\mathcal{C}\approx 1-\frac{3(\sigma_{2}^{2}+\sigma_{1}^{2}-2\sigma_{\text{he3}}^{2})}{4R^{2}(p_{T})}\approx 1-\frac{3\sigma_{2}^{2}}{4R^{2}(p_{T})},$ (8)
@@
which makes the ratio in Eq. (7) less than one and thus a large quantum correction. With $R(p_{T})$ decreasing with increasing $p_{T}$, the correction factor and thus the ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ ratio decreases with increasing $p_{T}$ as shown in Fig. 2 for all collision centralities.
For a small emission source with $R(p_{T})<\sigma_{1},\sigma_{\rm he3}\ll\sigma_{2}$, the correction factor becomes approximately
$\mathcal{C}\approx\frac{\sigma_{\text{he3}}^{6}}{\sigma_{1}^{3}\sigma_{2}^{3}}\left[1+3R^{2}(p_{T})\left(\frac{2}{\sigma_{\text{he3}}^{2}}-\frac{1}{\sigma_{1}^{2}}-\frac{1}{\sigma_{2}^{2}}\right)\right]$
$~{}\approx\frac{\sigma_{\text{he3}}^{6}}{\sigma_{1}^{3}\sigma_{2}^{3}}\left[1+\frac{3R^{2}(p_{T})}{\sigma_{\text{he3}}^{2}}\left(2-\frac{\sigma_{\text{he3}}^{2}}{\sigma_{1}^{2}}\right)\right],$ (9)
which also shows a suppression at large $p_{T}$ since the value of $R(p_{T})$ decreases as $p_{T}$ increases as a result of radial flow Aamodt et al. (2011); Acharya et al. (2024).
Suppression of hypertriton mean transverse momentum Shown in Fig. 3 by shaded bands is the ratio $\langle p_{T}\rangle_{{}^{3}_{\Lambda}\text{H}}/\langle p_{T}\rangle_{{}^{3}\text{He}}$ of hypertriton to helium-3 mean-$p_{T}$ as a function of the $\Lambda-d$ distance, with the solid symbols corresponding to results obtained with various values of $B_{\Lambda}$. This mean transverse momentum ratio is seen to be less than one and decreases as the $\Lambda-d$ distance increases or the hypertriton binding energy $B_{\Lambda}$ decreases with a variation on the order of 10%. This variation is much smaller than the factor of 2-3 variation for the yield ratio in Fig. 1 and the spectrum ratio in Fig. 2 because of the weak dependence of the ${}^{3}_{\Lambda}\text{H}/^{3}\text{He}$ ratio on the transverse momentum. Also shown by solid lines is the prediction of the blast-wave model, which indicates that the value of the hypertriton to helium-3 mean-$p_{T}$ ratio has a slightly larger than unity value of about 1.03 for the 0-10% collision centrality and about 1.04 for the 30-50% collision centrality. The different hypertriton to helium-3 mean-$p_{T}$ ratio in the coalescence model and the blast-wave model is unlike the suppressed hypertriton production in peripheral collisions, which can be described by the coalescence model with the hypertriton wave function and also by the thermal model with a canonical correlation volume Vovchenko et al. (2018). Therefore, the observation of the softening of hypertriton transverse momentum spectrum and the reduction of its mean transverse momentum, albeit their small magnitude, is of great significance because it clearly contradicts to the thermal model predictions and can serve as a clean probe to the hypertriton production mechanism in relativistic heavy-ion collisions.
Although a smaller mean $p_{T}$ for hypertriton than that for helium-3 is seen for all collision centralities, the centrality dependence of the softening effect shown in Fig. 3 is, however, rather weak because of the slowly decreasing radial flow with increasing collision centrality. Since the radial flow is even smaller and the $p_{T}$ dependence of $R(p_{T})$ is less prominent in collisions of small systems Aamodt et al. (2011); Adam et al. (2015b); Acharya et al. (2024), the softening of hypertriton transverse momentum spectrum in $pp$ and $pA$ collisions is expected to be even weaker in these collisions. For a quantitative study of the system size effect on the hypertriton transverse momentum spectrum requires, however, the extension of the present study to $pp$ and $pA$ collisions, which we leave for a future study.
@@
Conclusion and outlook. In the present study, we have investigated hypertriton production in Pb+Pb collisions at $\sqrt{s_{NN}}=5.02$ TeV using the coalescence model with kinetic freeze-out nucleons and $\Lambda$ hyperons from a microscopic hybrid approach based on the MUSIC hydrodynamic model and the UrQMD hadronic transport model. We have found that the halo structure of hypertriton with a large $\Lambda-d$ distance of approximately $10$ fm leads to not only a suppression of ${}^{3}_{\Lambda}\text{H}$ yield but also a softening of its transverse momentum ($p_{T}$) spectrum with a weak centrality dependence. In particular, the mean $p_{T}$ of ${}^{3}_{\Lambda}\text{H}$ is found to be smaller than that of helium-3 even in the most central collisions, which is in sharp contrast with the predictions of the blast-wave model. Such a quantum mechanical softening of (anti-)hypertriton spectrum is a general feature and a natural outcome of the hypertriton wave function used in the final-state coalescence model, which can be readily tested in high-energy experiments with different beam energies and collision systems, providing thus the possibility to unravel the production mechanism of (anti-)hypernuclei in high-energy nuclear collisions and to also obtain constraints on the $\Lambda$-nucleon interaction.